# LLMs as verification oracles for Solidity

### v1

@@ Abstract
Ensuring the correctness of smart contracts is critical, as even subtle flaws can lead to severe financial losses. While bug detection tools able to spot common vulnerability patterns can serve as a first line of defense, most real-world exploits and losses stem from errors in the contract business logic. Formal verification tools such as SolCMC and the Certora Prover address this challenge, but their impact remains limited by steep learning curves and restricted specification languages. Recent works have begun to explore the use of large language models (LLMs) for security-related tasks such as vulnerability detection and test generation. Yet, a fundamental question remains open: can LLMs serve as verification oracles, capable of reasoning about arbitrary contract-specific properties? In this paper, we provide the first systematic evaluation of GPT-5, a state-of-the-art reasoning LLM, in this role. We benchmark its performance on a large dataset of verification tasks, compare its outputs against those of established formal verification tools, and assess its practical effectiveness in real-world auditing scenarios. Our study combines quantitative metrics with qualitative analysis, and shows that recent reasoning-oriented LLMs can be surprisingly effective as verification oracles, suggesting a new frontier in the convergence of AI and formal methods for secure smart contract development and auditing.
@@ 1 Introduction
Recently, there has been a huge rise in works that adopt Large Language Models (LLMs) [11] for a variety of tasks in smart contract analysis, including common vulnerabilities detection [17, 18, 30, 47], test generation [6], and specification generation [28, 31]. While LLMs have been shown to perform reasonably well on such tasks, it remains unexplored whether they can also verify custom, contract-specific properties. This challenge is considerably more complex, since the properties to verify can be arbitrary and are strictly tied to the specific contract under analysis. Therefore, mere pattern recognition is not sufficient; a certain degree of reasoning capabilities is required. Recent models such as GPT-5 (released on August 7th, 2025) have claimed to show human-like reasoning capabilities; however, whether these capabilities transfer to smart contract verification has not been explored yet.
@@
Overall, our analysis shows that recent models such as GPT-5 can be effectively used to support smart contract auditing. To foster reproducibility, we release our dataset, the scripts needed to run the experiments, and the complete experimental results in a public (anonymised) repository [2].
@@ 2.1 Symbolic verification tools for Solidity
Both tools address the state explosion problem inherent in symbolic verification by approximating the Solidity semantics (e.g., both of them abstract from gas costs). These approximations, however, may yield false negatives (e.g., a property holds but the tool classifies it as violated), or false positives (the property is violated, but the tool reports it as satisfied). Assessing the reliability of a classification then requires a deep expertise on both the Solidity semantics and the specific approximations made by the tool, which in practice leads to a costly, iterative process of refining the specifications.
@@ 3 Dataset design
    LendingProtocol: this is the most complex contract in our controlled dataset, as it implements a subset of the functionality of the Aave protocol [1] (with some simplifications in order to keep the ground truth reliable). Its allow us to further explore non-linear arithmetic, integer truncations, and economic properties (e.g., certain actions produce certain gains for the involved users).
@@ 4 Experimental setup for LLMs evaluation
To mitigate bias toward simpler tasks, we did not feed the models with the entire dataset. Rather, we sampled verification tasks from the three simpler use cases, while retaining the complete set of tasks from the others. The resulting sub-dataset (hereafter, referred to as $\mathrm{DS}$) contains 667 verification tasks, and it is nearly balanced, featuring $\sim\mbox{1:1}$ ratio between simpler and complex use cases, and a comparable $\sim\mbox{1:1}$ ratio between positive and negative tasks.
@@ 5 RQ1: LLMs metrics on the controlled dataset
OVERALL (GPT-5) 667 0 302 309 32 24 92% 91% 93% 91% 92%
Table 2 reports the experimental results of GPT-5 and GPT-4 on dataset $\mathrm{DS}$. GPT-5 achieves an F1 score above $86\%$ in all use cases ($92\%$ overall), outperforming GPT-4 on all of them with a performance gap of $25\%-30\%$ on every prediction metric. This result aligns with prior findings on GPT-4’s limited logical reasoning abilities [29]. The most challenging use case was PaymentSplitter, where GPT-4 performs only slightly better than random guessing, while GPT-5 still maintains performances above 80% across all metrics. We observe that GPT-5 has a very low rate of FP, which is a desirable soundness property of verification tools. We elaborate in Section 6 the main causes of FPs and FNs.
@@ 6 RQ2: Analysis of LLMs explanations
To compare GPT-4 vs. GPT-5, we focussed on the Bank contract, since its simplicity allowed us to exhaustively evaluate all the answers. As we can see, coherence is rarely an issue, suggesting that both models are rather able to correctly interpret the property; this is further supported by the fact that slight variations of the properties (e.g., different assumptions) often lead to the expected explanations and counterexamples. The main cause of false results for GPT-4 is incorrectness. GPT-5, on the contrary, performs reasonably well on all criteria. In case of wrong answers, the cause distributes rather equally among the three criteria; in case of correct answers, the main weak point is completeness.
In order to answer to RQ2b, we analyzed the causes of incorrect answers. By manually inspecting the verification tasks where FPs are produced, we noted that they have a few common causes: 1 low-level calls triggering ETH transfers; 2 selfdestructtransactions increasing the contract balance; 3 integer overflows (mostly, in PaymentSplitter). We note, however, that GPT-5 is not consistently wrong when dealing with these features: rather, it answers correctly most of the times. A typical cause for FNs is the use of transfer to send ETH: here, GPT-5 often misses that this operation does not carry enough ETH to allow the callee to perform a selfdestruct.
@@
Use case # SolCMC Certora GPT-4 GPT-5
Bank 121 62.81   (61.98) 89.26   (89.26) 100   (100) 100   (100)
Vault 106 26.42   (26.42) 54.72   (54.72) 100   (96.23) 100   (100)
PriceBet 158 47.47   (46.84) 56.33   (56.33) 100   (97.47) 100   (100)
PaymentSplitter 160 43.75   (18.13) 87.50   (87.50) 100   (98.75) 100   (100)
LendingProtocol 122 4.10   (0) 81.15   (77.05) 100   (99.18) 100   (100)
OVERALL 667 38.08   (30.89) 74.06   (73.31) 100   (98.35) 100   (100)
As discussed in Section 3, only certain properties can be expressed in the specification languages of SolCMC and Certora. We define expressibility as the percentage of verification tasks that can be expressed in the given tool (for LLMs, this is, by default, $100\%$). We also measure the coverage, i.e. the percentage of tasks for which the tool returns True or False (instead of Unknown or timeouts).
@@
$\mathrm{\mathrm{DS}_{|CVL}}$ Certora 494 5 196 179 58 56 77% 77% 78% 76% 77%
$\mathrm{\mathrm{DS}_{|CVL}}$ GPT-5 494 0 238 214 24 18 91% 91% 93% 90% 92%
$\mathrm{\mathrm{DS}_{|\neg CVL}}$ GPT-5 173 0 63 96 9 5 92% 88% 93% 91% 90%
$\mathrm{\mathrm{DS}_{|CVL}}$ GPT-4 494 6 173 149 86 80 66% 67% 68% 63% 68%
$\mathrm{\mathrm{DS}_{|\neg CVL}}$ GPT-4 173 5 37 55 46 30 55% 45% 55% 54% 49%
@@
$\mathrm{\mathrm{DS}_{|\neg SolCMC}}$ GPT-5 413 0 191 191 14 17 92% 93% 92% 93% 92%
@@
Besides correctly identifying the validity of a property, the value of a verification tool lies in the guarantees it can provide about the answer (i.e., a counterexample or proof certificate). In Section 6, we analyzed the answers of GPT-4 and GPT-5. Here, we discuss how they compare with those of SolCMC and Certora. Both tools, when returning True, do not provide any explanation: the guarantees just lie on their correct implementation and on their symbolic interpretation of the Solidity semantics (but, as observed before, some approximations they make are unsound). When returning False, SolCMC tries to return a sequence of transactions that brings the contract in a state where the invariant is violated. In most experiments, however, it did not produce any counterexample; moreover, it may return FNs in certain cases in which it abstracts from the Solidity semantics [42]. Certora, on the contrary, over-approximates the set of reachable states: its counterexamples are states which violate the property, but they are not guaranteed to be reachable, thereby producing spurious counterexamples [13]. This explains the high number of FNs in Table 5.
@@ 8 RQ4: LLMs for verification auditing in the wild
The reports contain properties written in natural language, each of which is associated to a formal specification written in CVL. We feed GPT-5 the properties written in natural language, and ask it to verify them with the same prompt used in the other experiments (i.e., the one in Appendix 0.A). We stress that, since GPT-5 knowledge cut-off is September 30th 2024 [36], we can exclude that the model uses information found in the reports. This is further validated by the fact that GPT-5 often disagrees with Certora.
In general, we cannot determine the ground truth of the properties, as these contracts consist of thousands of lines of code, besides those in external dependencies. Hence, reliably establishing the validity of their properties is rather unfeasible, forcing us to recur to a more empirical approach.
We carefully analysed GPT-5 answers, with particular attention to properties over which GPT-5 disagrees with Certora. First, we examined the CVL specification, to verify whether it is consistent with the property written in natural language; if not, we rephrased the property accordingly, and re-ran GPT-5 over the newly formulated property. If the cause of disagreement is not an inconsistency between the formal specification and the property description, we then analysed the counterexample provided by one of the two tools. In particular, since the states that Certora provides as counterexamples lie on an over-approximation of the set of reachable states, we had to ensure that the returned state is indeed reachable, in order to confirm the violation. In order to do that, we paired manual inspection with new additional queries to GPT-5, feeding it the Certora’s state and asking it to provide a trace that starts from the deployment of the contract and leads to the state under analysis.
@@ 9 Related work
Research on smart contracts analysis has traditionally followed two main directions: vulnerability detection and formal verification. The former focusses on finding bugs in contract code by leveraging recurring vulnerability patterns. In this setting, there exist large datasets of contracts tagged with their vulnerabilities that are used as benchmarks to evaluate new analysis techniques [38, 41, 45]. Formal verification instead focusses on proving that a contract satisfies specific properties of its business logic (or showing counterexamples otherwise). This approach is pursued through different perspectives, often related to the class of expressible properties. For example, VeriSolid [35] focusses on CTL properties; SmartPulse [39] on past-LTL properties; Solvent [8] on strategic properties of the form “there exists some transaction leading to a certain state”. Compared to SolCMC and Certora, however, these tools are not yet directly usable in the wild, as they typically support restricted fragments of Solidity. Unlike the vulnerability detection setting, no large datasets exist for benchmarking formal verification tools. The largest available dataset so far contains only 323 verification tasks [9]; whereas the dataset developed for this work comprises 2034 tasks.
@@
LLMs offer a complementary path to contract analysis: thanks to their understanding of natural language and reasoning capabilities, they may help overcoming some of the limitations of formal verification tools. David et al. [17] explore the use of LLMs for smart contract auditing, by querying GPT-4 and Claude to detect the presence of a pre-defined vulnerability patterns (e.g. reentrancy, delegatecall injection, etc.). Using a benchmark of 52 compromised DeFi contracts, they show that GPT-4 and Claude accurately identify vulnerabilities in ${\sim}40\%$ of cases, albeit the high FP rate still requires human oversight. At the light of our results, we believe that GPT-5 would substantially improve the metrics of [17]. SmartGuard [18] addresses vulnerability detection by combining the retrieval of semantically similar code with chain-of-thought generation and a self-check architecture, validating their technique on the SolidiFI benchmark [19]. The work [30] fine-tunes ChatGPT on 7 logical vulnerabilities (e.g., price oracle manipulation, privilege escalation), and experiments on a dataset of audit reports. LLM-SmartAudit [47] employs a multi-agent conversational framework with a buffer-of-thought mechanism, addressing 10 types of common vulnerabilities (reentrancy, overflow, etc.). While the previous three works achieve high prediction metrics on their respective datasets, the use of different datasets for validation makes direct comparison challenging.
@@
Compared to the previous works, ours takes a different perspective: rather than focussing on detecting known vulnerability types, we investigate whether LLMs are capable of reasoning about arbitrary, contract-specific properties, and whether they can provide convincing explanations for their answers.
@@
During the construction of our dataset, we observed that crafting properties that precisely capture the desirable contract behaviors is quite challenging. To address this issue, some works have proposed LLMs as an aid to automate the generation of such properties. PropertyGPT [31] uses LLM-based in-context learning to generate properties from contract code, targeting in a custom specification language similar to Certora’s CVL. The work [28] also uses LLMs, targeting post-conditions in the language of solc-verify [22]. A limitation of both works is that they cannot deal with arbitrary properties, since after generation they still resort to symbolic tools for verification. As noted before, this rules out relevant classes of properties (e.g., strategic ones), which could provide more precise insights on the desirable contract dynamics. Indeed, it has been observed that “global” properties (such as metamorphic or strategic properties) usually yield a higher return of investment than “local” ones, such as function specs [48].
@@
While our work only hints at the potential of combining LLMs with formal methods, other studies — in domains outside of smart contracts — have already shown that such integration can be quite effective in practice. For example, ESBMC-AI [43] integrates LLMs with bounded model checking to detect and repair common flaws. The work [26] show that LLMs can aid in formal verification tasks, such as inferring loop invariants and proving safety properties.

### v2

@@ Abstract
Ensuring the correctness of smart contracts is critical, as even subtle flaws can lead to severe financial losses. While bug detection tools able to spot common vulnerability patterns can serve as a first line of defense, most real-world exploits and losses stem from errors in the contract business logic. Formal verification tools such as SolCMC and the Certora Prover address this challenge, but their impact remains limited by steep learning curves and restricted specification languages. Recent works have begun to explore the use of large language models (LLMs) for security-related tasks such as vulnerability detection and test generation. Yet, a fundamental question remains open: can LLMs aid in assessing the validity of arbitrary contract-specific properties? In this paper, we provide the first systematic empirical evaluation of GPT-5, a state-of-the-art reasoning LLM, in this role. We benchmark its performance on a large dataset of verification tasks, compare its outputs against those of established formal verification tools, and assess its practical effectiveness in real-world auditing scenarios. Our study combines quantitative metrics with qualitative analysis, and shows that recent reasoning-oriented LLMs — although lacking soundness guarantees — can be surprisingly effective at predicting the (in)validity of complex properties, suggesting a new frontier in the convergence of AI and formal methods for secure smart contract development and auditing.
@@ 1 Introduction
Recently, there has been a huge rise in works that adopt Large Language Models (LLMs) [12] for a variety of tasks in smart contract analysis, including common vulnerabilities detection [19, 21, 35, 53], test generation [6], and specification generation [33, 36]. While LLMs have been shown to perform reasonably well on such tasks, it remains unexplored whether they can also play a role in the verification of custom, contract-specific properties. This challenge is considerably more complex, since the properties to verify can be arbitrary and are strictly tied to the specific contract under analysis. Therefore, mere pattern recognition is not sufficient; a certain degree of reasoning capabilities is required. Recent models such as GPT-5 (released on August 7th, 2025) have claimed to show human-like reasoning capabilities; however, whether these capabilities transfer to smart contract verification has not been explored yet.
@@
Overall, our analysis shows that recent models such as GPT-5 can be effectively used to support smart contract auditing. To foster reproducibility, we release our dataset, the scripts needed to run the experiments, and the complete experimental results in a public repository [43].
@@ 2.1 Symbolic verification tools for Solidity
Both tools make some approximations of the Solidity semantics (e.g., both of them abstract from gas costs) to make verification more practical. These approximations, however, may yield false negatives (e.g., a property holds but the tool classifies it as violated), or false positives (the property is violated, but the tool reports it as satisfied). Assessing the reliability of a classification then requires a deep expertise on both the Solidity semantics and the specific approximations made by the tool, which in practice leads to a costly, iterative process of refining the specifications.
@@ 3 Dataset design
    LendingProtocol: this is the most complex contract in our controlled dataset, as it implements a subset of the functionality of the Aave protocol [1] (with some simplifications in order to keep the ground truth reliable). It allows us to explore non-linear arithmetic, integer truncations, and economic properties (e.g., certain actions produce certain gains for the involved users[10]).
@@ 4 Experimental setup for LLMs evaluation
To mitigate bias toward simpler tasks, we did not feed the models with the entire dataset. Rather, we sampled verification tasks from the three simpler use cases, while retaining the complete set of tasks from the others. The resulting sub-dataset (hereafter, referred to as $\mathrm{DS}$) contains 667 verification tasks, and it is nearly balanced, featuring $\sim\mbox{1:1}$ ratio between simpler and complex use cases, and a comparable $\sim\mbox{1:1}$ ratio between positive and negative tasks. Text generation used the API’s default decoding parameters at the time of the experiments. Specifically, the sampling temperature was set to 1.0, nucleus sampling was disabled (top-p = 1.0), and no frequency or presence penalties were applied. The maximum number of generated tokens was 30,000, with generation terminating either upon reaching this limit or according to the model’s internal end-of-sequence criteria. The total cost of the experiments was below €100. The full dataset—including scripts, formal specifications in SolCMC and Certora, and all LLM outputs, is available in [43].
@@ 5 RQ1: LLMs metrics on the controlled dataset
OVERALL (GPT-5) 667 0 303 312 31 21 92% 91% 94% 91% 92%
Table 2 reports the experimental results of GPT-5 and GPT-4 on dataset $\mathrm{DS}$. GPT-5 achieves an F1 score above $88\%$ in all use cases ($92\%$ overall), outperforming GPT-4 on all of them with a performance gap of $25\%-30\%$ on every prediction metric. This result aligns with prior findings on GPT-4’s limited logical reasoning abilities [34]. The most challenging use case was PaymentSplitter, where GPT-4 performs only slightly better than random guessing, while GPT-5 still maintains performances above 80% across all metrics. We observe that GPT-5 has a very low rate of FP, which is a desirable soundness property of verification tools. We elaborate in Section 6 the main causes of FPs and FNs.
@@ 6 RQ2: Analysis of LLMs explanations
To compare GPT-4 vs. GPT-5, we focussed on the Bank contract, on which we exhaustively evaluated all the answers. Coherence is rarely an issue, suggesting that both models are able to correctly interpret the property; this is further supported by the fact that slight variations of the properties (e.g., different assumptions) often lead to the expected explanations and counterexamples. The main cause of false results for GPT-4 is incorrectness. GPT-5, on the contrary, performs reasonably well on all criteria. For wrong answers, the cause distributes rather equally among the three criteria; for correct answers, the main weak point is completeness. Since the simplest use case already shows a significant gap in reasoning capabilities between the two models, we did not extend the qualitative comparison of GPT-4’s and GPT5’s answers to the other use cases.
To address RQ2b, we analyzed the causes of incorrect answers. By manually inspecting the verification tasks that produced FPs, we identified a few recurring issues: 1 low-level calls triggering ETH transfers; 2 selfdestructtransactions that increase a contract’s balance; 3 integer overflows (primarily, in PaymentSplitter). We note, however, that GPT-5 is not consistently incorrect on these features, as it produces correct answers in most cases. A common cause of FNs is the use of transfer to send ETH: here, GPT-5 often misses that this operation does not carry enough gas to allow the callee to perform a selfdestruct.
@@
Use case # SolCMC Certora
Bank 121 62.81   (61.98) 89.26   (89.26)
Vault 106 26.42   (26.42) 54.72   (54.72)
PriceBet 158 47.47   (46.84) 56.33   (56.33)
PaymentSplitter 160 43.75   (18.13) 87.50   (87.50)
LendingProtocol 122 4.10   (0) 81.15   (77.05)
OVERALL 667 38.08   (30.89) 74.06   (73.31)
As discussed in Section 3, only certain properties can be expressed in the specification languages of SolCMC and Certora. We define expressibility as the percentage of verification tasks that can be expressed in the given tool. We also measure the coverage, i.e. the percentage of tasks for which the tool returns True or False (instead of Unknown or timeouts).
@@
$\mathrm{\mathrm{DS}_{|CVL}}$ Certora 494 5 196 178 58 57 76% 77% 77% 75% 77%
$\mathrm{\mathrm{DS}_{|CVL}}$ GPT-5 494 0 240 215 22 17 92% 92% 93% 91% 92%
$\mathrm{\mathrm{DS}_{|\neg CVL}}$ GPT-5 173 0 63 97 9 4 92% 88% 94% 92% 91%
$\mathrm{\mathrm{DS}_{|CVL}}$ GPT-4 494 6 174 149 85 80 66% 67% 69% 64% 68%
$\mathrm{\mathrm{DS}_{|\neg CVL}}$ GPT-4 173 5 36 55 47 30 54% 43% 55% 54% 48%
@@
$\mathrm{\mathrm{DS}_{|\neg SolCMC}}$ GPT-5 413 0 193 193 12 15 93% 94% 93% 94% 93%
@@
Besides correctly identifying the validity of a property, the value of a verification tool lies in the guarantees it can provide about the answer (i.e., a counterexample or proof certificate). In Section 6, we analyzed the answers of GPT-4 and GPT-5. Here, we discuss how they compare with those of SolCMC and Certora. Both tools, when returning True, do not provide any explanation: the guarantees just lie on their correct implementation and on their symbolic interpretation of the Solidity semantics (but, as observed before, some approximations they make are unsound). When returning False, SolCMC tries to return a sequence of transactions that leads the contract in a state where the invariant is violated. In most experiments, however, it did not produce any counterexample; moreover, it may return FNs in certain cases in which it abstracts from the Solidity semantics [48]. Certora, on the contrary, over-approximates the set of reachable states: its counterexamples are states that violate the property, but they are not guaranteed to be reachable, thereby producing spurious counterexamples [14]. This explains the high number of FNs in Table 5.
@@ 8 RQ4: LLMs for verification auditing in the wild
The reports contain properties written in natural language, each of which is associated to a formal specification written in CVL. We feed GPT-5 the properties written in natural language, and ask it to verify them with the same prompt used in the other experiments (i.e., the one in Appendix 0.A). We stress that, since GPT-5 knowledge cut-off is September 30th 2024 [41], we can assume that the model does not use information found in the reports. This is further validated by the fact that GPT-5 often disagrees with Certora.
In general, determining the ground truth of the properties is extremely challenging, as these contracts consist of thousands of lines of code, besides those in external dependencies. Hence, reliably establishing the validity of all the properties is rather unfeasible, forcing us to recur to a more empirical approach.
We carefully analysed GPT-5 answers, with particular attention to properties over which GPT-5 disagrees with Certora. First, we examined the CVL specification, to verify whether it is consistent with the property written in natural language; if not, we rephrased the property accordingly, and re-ran GPT-5 over the newly formulated property. If the cause of disagreement is not an inconsistency between the formal specification and the property description, we then analysed the counterexample provided by one of the two tools. In particular, since the states that Certora provides as counterexamples lie on an over-approximation of the set of reachable states, we had to ensure that the returned state is indeed reachable, in order to confirm the violation. In order to do that, we paired manual inspection with new additional queries to GPT-5, feeding it the Certora’s state and asking it to provide a trace that starts from the deployment of the contract and leads to the state under analysis. For a sample of the properties deemed as False by GPT-5 but True by Certora, we produced Proofs of Concept (available in [43]) written in Forge [22] that validate GPT-5’s answers.
@@ 9 Related work
Research on smart contracts analysis has mainly followed two directions: vulnerability detection and formal verification. The former focusses on finding bugs in contract code by leveraging recurring vulnerability patterns. In this setting, there exist large datasets of contracts tagged with their vulnerabilities that are used as benchmarks to evaluate new analysis techniques [44, 47, 51]. Formal verification instead focusses on proving that a contract satisfies specific properties of its business logic (or showing counterexamples otherwise). This approach is pursued through different perspectives, often related to the class of expressible properties. For example, VeriSolid [40] focusses on CTL properties; SmartPulse [45] on past-LTL properties; Solvent [8] on strategic properties of the form “there exists some transaction leading to a certain state”. Compared to SolCMC and Certora, however, these tools are not yet directly usable in the wild, as they typically support restricted fragments of Solidity. Unlike the vulnerability detection setting, no large datasets exist for benchmarking formal verification tools. The largest available dataset so far contains only 323 verification tasks [9]; whereas the dataset developed for this work comprises 2034 tasks.
@@
LLMs offer a complementary path to contract analysis: thanks to their understanding of natural language and pattern-driven problem-solving approach, they may help overcoming some of the limitations of formal verification tools. A1 [23] and ReX [54] leverage LLMs — paired with testing and concrete execution — to generate exploit against on-chain contracts, showing that they are are able to reproduce exploits for several historical vulnerabilities. The role of LLMs in the context of smart contract auditing has been first analysed [19], where GPT-4 and Claude are queried to detect the presence of a pre-defined vulnerability patterns (e.g. reentrancy, delegatecall injection, etc.). SmartGuard [21] addresses vulnerability detection by combining the retrieval of semantically similar code with chain-of-thought generation and a self-check architecture, validating their technique on the SolidiFI benchmark [24]. The work [35] fine-tunes ChatGPT on 7 logical vulnerabilities (e.g., price oracle manipulation, privilege escalation), and experiments on a dataset of audit reports. LLM-SmartAudit [53] employs a multi-agent conversational framework with a buffer-of-thought mechanism, addressing 10 types of common vulnerabilities (reentrancy, overflow, etc.).
@@
Compared to the previous works, ours takes a different perspective: rather than focussing on detecting known vulnerability types, and possibly use them to synthesize exploits, we investigate whether LLMs are capable of reasoning about arbitrary, contract-specific properties, and whether they can provide convincing explanations for their answers.
@@
During the construction of our dataset, we observed that crafting properties that precisely capture the desirable contract behaviors is quite challenging. To address this issue, some works have proposed LLMs as an aid to automate the generation of such properties. PropertyGPT [36] uses LLM-based in-context learning to generate properties from contract code, targeting in a custom specification language similar to Certora’s CVL. The framework DbC-GPT [33] proposes to use LLMs to generate post-conditions in the language of solc-verify [27]. A limitation of both works is that they cannot deal with arbitrary properties, since after generation they still resort to symbolic tools for verification. As noted before, this rules out relevant classes of properties (e.g., strategic ones), which could provide more precise insights on the desirable contract dynamics. Indeed, it has been observed that “global” properties (such as metamorphic or strategic properties) usually yield a higher return of investment than “local” ones, such as function specs [55].
@@
While our work only hints at the potential of combining LLMs with formal methods, other studies — in domains outside of smart contracts — have already shown that such integration can be quite effective in practice. For example, ESBMC-AI [49] integrates LLMs with bounded model checking to detect and repair common flaws. The work [31] shows that LLMs can aid in formal verification tasks, such as inferring loop invariants and proving safety properties.
10 Limitations
Soundness.
The use of LLMs in the verification context cannot replace — nor it should be intended as a substitute for — traditional formal verification approaches. Indeed, LLMs’ predictions and explanations — although statistically accurate — are not supported by sound reasoning and are not certifiable. Nevertheless, both the experiments on the synthetic dataset and those on real-world reports suggest that LLMs can be profitably used in combination with symbolic techniques, e.g. concrete PoCs that refine and certify the LLM-produced counterexamples.
Limitations of the dataset.
Although our dataset includes a wide range of property classes — including state invariants, single- and multi-transition invariants, function specs, metamorphic and strategic properties — it does not encompass all relevant classes of properties. In particular, some contracts, such as gambling games and certain DeFi protocols, rely on incentive mechanisms whose correctness can only be analysed in a game-theoretic setting. For example, in an $N$-players lottery, one would be interested in proving fairness properties, such as that each honest player will have a non-negative payoff [4, 39, 18]. Additionally, DeFi protocols are usually designed to achieve complex economic properties (e.g., “market efficiency” for AMMs). These kinds of properties are beyond the scope of existing industrial verification tools such as SolCMC and Certora. Investigating whether LLMs can provide reliable predictions and explanations for these classes of properties is a largely unexplored direction for future work.
Additional limitations of our dataset concern its scale, the subset of Solidity features it explores, and its coverage of real-world contract patterns. In particular, the contract variants included in the dataset do not expose some potentially problematic constructs such as the gas mechanism, loops, and delegate calls. Moreover, the included contracts do not cover some patterns used in practice, such as proxies, factories, and cross-chain interactions [56]. Assessing the LLMs’ understanding of these features is another direction for future work.